Solve for x and show your work
3x-2(-50/19)=40

Answer by stanbon(75887)

Solve for x and show your work
3x-2(-50/19)=40
---------
3x +100/19 = 40
------
57x + 100 = 760
---
57x = 660
x = 11.58
=============================
Cheers,
Stan H.